OpenStudy (anonymous):

Phil Maker is buying a camera for $100, including 6% sales tax. Find the cost of the camera before sales tax.

OpenStudy (anonymous):

x+.6x=100

OpenStudy (anonymous):

1.6x=100

OpenStudy (anonymous):

x=100/1.6

OpenStudy (anonymous):

x=62.5

OpenStudy (anonymous):

it cost $62.50

OpenStudy (anonymous):

let me check my work because that doesn't seem right...

OpenStudy (anonymous):

Thank you but actually the answer is $94.34

OpenStudy (anonymous):

where did I goof up?

OpenStudy (anonymous):

100=x+0.06x 100= 1.06x 100/1.06= x 94.33962264 = x round = $94.34

OpenStudy (anonymous):

NANCY, THANK YOU SO MUCH FOR SHOWING THE CORRECT WAY TO FIGURE OUT TH PROBLEM.

OpenStudy (anonymous):

GRANDALFWIZ, THANK YOU AS WELL. YOU STARTED IT OFF BUT WE DID NOT WORK IT OUT

OpenStudy (anonymous):

right.... .06, not .6 sorry!

OpenStudy (anonymous):

It is ok, i appreciate the help